OpenStudy (anonymous):

graph y+3=-2(x+4) which is y=-2x+5 correct

OpenStudy (anonymous):

incorrect, y+3=-2(x+4) is y=-2x-11

OpenStudy (anonymous):

make sure to distribute the negative sign when you multiply 4 by negative 2

OpenStudy (anonymous):

make sure to SUBTRACT 3 from both sides (don't add 3)